If 0.00100 moles of hydrogen gas at 155 C exert a pressure of 323 mmHg. What is the volume?


1
Expert's answer
2021-05-20T07:54:16-0400

PV=nRT

R= 8.314

T= 273+155=428


"V=\frac{nRT}{P}"


"V=\frac{0.00100×8.314×428}{323}=0.011"